urfavvamanii:

what's 2 +2 help a sista outπŸ˜”

Stringcheese272:

bruh

Illuminous:

22 duh.

Stringcheese272:

@illuminous wrote:
22 duh.
fr

urfavvamanii:

noo i think it'd 44

5172319:

its 11 yall for anyone who didnt know

urfavvamanii:

nah 2+2 equals 90

urfavvamanii:

11?

Illuminous:

it's 22!

urfavvamanii:

ITS 90

Illuminous:

No no all wrong..

5172319:

yes its defitly 11

Illuminous:

It's 21

urfavvamanii:

su b4 i cake u down

Illuminous:

TWENTY ONE

5172319:

@urfavvamanii wrote:
su b4 i cake u down
woahhhh

urfavvamanii:

@5172319 wrote:
@urfavvamanii wrote:
su b4 i cake u down
woahhhh
shh

Illuminous:

TWENTY ONEEE CAN U DO SUM FOR MEEEEEEEE

urfavvamanii:

@illuminous wrote:
TWENTY ONEEE CAN U DO SUM FOR MEEEEEEEE
yes lemme dic u down

urfavvamanii:

we gon have baby

urfavvamanii:

I give u baby

5172319:

@urfavvamanii wrote:
@5172319 wrote:
@urfavvamanii wrote:
su b4 i cake u down
woahhhh
shh
😏 😏

urfavvamanii:

πŸ˜›

urfavvamanii:

@5172319 wrote:
@urfavvamanii wrote:
@5172319 wrote:
@urfavvamanii wrote:
su b4 i cake u down
woahhhh
shh
😏 😏
bye

5172319:

@urfavvamanii wrote:
@5172319 wrote:
@urfavvamanii wrote:
@5172319 wrote:
@urfavvamanii wrote:
su b4 i cake u down
woahhhh
shh
😏 😏
bye
goodbye

urfavvamanii:

okOK

R3B3LMARINE:

2+2= me and you

Illuminous:

hey LILLY

5172319:

@r3b3lmarine wrote:
2+2= me and you
REALLLLL

urfavvamanii:

yall are gay

5172319:

@urfavvamanii wrote:
yall are gay
and so are u

5172319:

but i dont support that behavior

urfavvamanii:

@5172319 wrote:
@urfavvamanii wrote:
yall are gay
and so are u
gay 4 lilyπŸ˜›

Illuminous:

LILLY

Illuminous:

theres two lillys on this post

urfavvamanii:

oh

5172319:

@illuminous wrote:
LILLY
Im just confused atpp bc im lily and ur lily

urfavvamanii:

I;m tb the black onee

urfavvamanii:

ARE U BOTH BLACK

5172319:

@illuminous wrote:
@5172319 wrote:
@illuminous wrote:
LILLY
Im just confused atpp bc im lily and ur lily
Ik bbg mwah
mwah

Illuminous:

NO

Illuminous:

SHES WHITE

urfavvamanii:

OH

urfavvamanii:

OHH

5172319:

no im white

Illuminous:

YESSSSS SHES WHITEEEEE

urfavvamanii:

HI LKIKLUY

urfavvamanii:

πŸ˜›

5172319:

hold on wait a minute

5172319:

now im confused

urfavvamanii:

I got faith in u nd I

urfavvamanii:

BABBBBBBBBAAAAAAA GRILL

5172319:

@urfavvamanii wrote:
HI LKIKLUY
to who me?

Midnight97:

2+2=4 that's what i mean

urfavvamanii:

@5172319 wrote:
@urfavvamanii wrote:
HI LKIKLUY
to who me?
are u lily?

smith123:

21?

5172319:

@urfavvamanii wrote:
@5172319 wrote:
@urfavvamanii wrote:
HI LKIKLUY
to who me?
are u lily?
yes